OpenStudy (anonymous):

-1/2 (-2,-3) in standard form?

OpenStudy (anonymous):

standard form of what?

OpenStudy (anonymous):

line?

OpenStudy (anonymous):

as in Ax+By=C

OpenStudy (anonymous):

its a line n i guess u r given slope n point on the line,

OpenStudy (anonymous):

idk thats all the question gives me

OpenStudy (anonymous):

y+3=-1/2(x+2) y+3=-1/2x-1 1/2x+y=-1-3 1/2x+y=-4